### v1

@@ Appendix B IP Certificates in Restricted Networks
The simple form of retrieving temporal IP certificates is not feasible if the network operator does not tolerate the usage of Seldom, e.g., how the Chinese Firewall blocks regular Tor traffic (Dunna et al., 2018). Certificate requests are easily interceptable and blockable, or worst, the requester could be prosecuted. Therefore, we discuss in the following how someone who wishes to use Seldom could securely request IP certificates under these circumstances. In the past, much effort was invested into making Tor traffic opaque to traffic analysis. Similarly to how bridge relays works, CAs and clients can communicate in a way that does not resemble classical certificate requests. For this, either the CA needs access to enough different IP addresses that not all of them can be blocked simultaneously, or the CA can accept certificate requests through proxies. The process to request certificates in the latter case works as follows. Because the CA cannot verify the origin of a request, the host of the proxy server could request certificates for whatever IP address he desires. Therefore, the certificate itself is not distributed over the same proxy as the request. Instead, the proxy is only informed about other servers that will host shares of the certificate and communicates this to the requesting client. The client can then establish covert connections to these servers and retrieves shares of its certificate. Each of these new proxies verifies the IP address of the client, which ensures that certificates are only created for the user controlling the certified IP address. Only a single honest proxy is necessary to ensure correctness and also identify misbehavior of other proxies.
@@ Appendix C Hybrid Threshold Encryption
For Seldom, we designed and implemented a hybrid threshold-encryption scheme based on RSA and AES which is published as the thRSAhold Python libary.111thRSAhold is available on PyPi and on GitHub at https://github.com/eric-wagner/thRSAhold. Our thRSAhold library is compatible with standard RSA and AES implemenations, such that we could use openssl for fast encryption and relied only on slower Python code for deanonymizations.
Our threshold-encryption closely follows Shoup’s threshold RSA signatures (Shoup, 2000). All plaintexts are padded, to be at least as long as the modulus of the public RSA key to prevent common attacks against RSA. If the plaintext does not fit into a single RSA ciphertext, a random AES key is generated and prepended to the plaintext. This key and the beginning of the original plaintext message are threshold-encrypted, while the remaining plaintext is encrypted with AES in GCM mode to ensure condifentiality and data authenticity.
During decryption, the first part of the message, clearly defined by the size of the public RSA modulus, is threshold decrypted by the consortium. Only an entity that collects enough decryption shares can then reconstruct the first part of the plaintext, which only requires knowledge about the public encryption key. Thus, the requesting LEA can decrypt the identity of the suspicious user in the final deanonymization step, without revealing this information to the consortium itself. If the ciphertext is longer than the RSA modulus, the first of the decrypted plaintext are interpreted as AES key to decrypt the remaining ciphertext. The key can then be discarded afterwards. Finally, the padding is removed from the plaintext to reveal the original plaintext.
@@ Appendix D Consortium Smart Contract
Listing 1 shows an examplary Solidity smart contract to manage a static consortium of five members with a basic majority vote to accept a deanonymization request. The highlighed lines of code show the variables and events that are revealed to the involved exit node to prove that the deanonymization request for a specific flow record has been accepted and to the public (either in real time or through delayed disclosure).

### v2

@@ Appendix B Opaquely Retrieving IP Certificates
The simple form of retrieving temporal IP certificates is not feasible if the network operator does not tolerate the usage of Seldom, e.g., how the Chinese Firewall blocks regular Tor traffic (Dunna et al., 2018). Certificate requests are easily interceptable and blockable, or worse, the requester could be prosecuted. Therefore, we discuss in the following how someone who wishes to use Seldom could securely request IP certificates under these circumstances. In the past, much effort was invested into making Tor traffic opaque to traffic analysis. Similarly to how bridge relays work, CAs and clients can communicate in a way that does not resemble classical certificate requests. For this, either the CA needs access to enough different IP addresses that not all of them can be blocked simultaneously, or the CA can accept certificate requests through proxies. The process to request certificates in the latter case works as follows. Because the CA cannot verify the origin of a request, the host of the proxy server could request certificates for whatever IP address they desire. Therefore, the certificate itself is not distributed over the same proxy as the request. Instead, the proxy is only informed about other servers that will host shares of the certificate and communicates this to the requesting client. The client can then establish covert connections to these servers and retrieve shares of its certificate. Each of these new proxies verifies the IP address of the client, which ensures that certificates are only created for the user controlling the certified IP address. A single honest proxy suffices to ensure correctness and also identify misbehavior of other proxies.
@@ Appendix C Hybrid Threshold Encryption
To efficiently implement Seldom, we designed and implemented a hybrid threshold-encryption scheme based on RSA and AES, which is published as the thRSAhold Python library.333thRSAhold is available on PyPi and on GitHub at https://github.com/eric-wagner/thRSAhold. Our thRSAhold library is compatible with standard RSA and AES implementations, such that, e.g., openssl can be used for fast encryption, and the slower Python code is only used for deanonymizations.
Our threshold encryption closely follows Shoup’s threshold RSA signatures (Shoup, 2000). All plaintexts are padded to be at least as long as the modulus of the public RSA key to prevent common attacks against RSA. If the plaintext does not fit into a single RSA ciphertext, a random AES key is generated and prepended to the plaintext. This key and the beginning of the original plaintext message are threshold-encrypted, while the remaining plaintext is encrypted with AES in GCM mode to ensure confidentiality and data authenticity.
During decryption, the first part of the message, clearly defined by the size of the public RSA modulus, is threshold decrypted by the consortium. Only an entity that collects enough decryption shares can then reconstruct the first part of the plaintext, which only requires knowledge about the public encryption key. Thus, the requesting LEA can decrypt the identity of the suspicious user in the final deanonymization step, without revealing this information to the consortium itself. If the ciphertext is longer than the RSA modulus, the first of the decrypted plaintext are interpreted as an AES key to decrypt the remaining ciphertext. The key can then be discarded afterward. Finally, the padding is removed from the plaintext to reveal the original plaintext.
@@ Appendix D Consortium Smart Contract
Listing 1 shows an exemplary Solidity smart contract to manage a static consortium of five members with a basic majority vote to accept a deanonymization request. The highlighted lines of code show the variables and events that are revealed to the involved exit node to prove that the deanonymization request for a specific flow record has been accepted and to the public (either in real time or through delayed disclosure).